OpenStudy (anonymous):

Which of these is a simplified form of the equation 7z + 2 = 8 + 2z + 1z? A)10z = 6 B)10z = 10 C)4z = 6 D)4z = 10

OpenStudy (anonymous):

7z + 2 = 8 + 2z + z 7z + 2 = 3z + 8 (Simplify) 4z + 2 = 8 (Take 3z from both sides) 4z = 6 (Take 2 from each side)

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thank you